# Dynamic Multi-Expert Aggregation

### v1

@@ Abstract
Shoubin Yuthanks: Equal contribution.  Yue Zhang  Ziyang Wang Jaehong YoonMohit Bansal UNC Chapel Hill https://github.com/Yui010206/MEXA
@@ 1 Introduction
The rapid advancement of multimodal learning has significantly improved AI systems’ ability to understand, reason, and interact with the real world, benefiting diverse tasks such as visual question answering Antol et al. (2015); Hudson and Manning (2019); Marino et al. (2019); Mathew et al. (2021); Tanaka et al. (2023); Winata et al. (2024); Wang et al. (2025); Zhang et al. (2024b), medical image diagnosis Lau et al. (2018); He et al. (2020); Liu et al. (2021); Azam et al. (2022); Cai et al. (2023); Bai et al. (2024), and embodied AI Brohan et al. (2023); Driess et al. (2023); Liu et al. (2024b); Durante et al. (2024); Majumdar et al. (2024); Zhang et al. (2024c). As AI systems become more integrated into complex, real-world applications, the ability to flexibly understand and reason upon heterogeneous multimodal inputs is increasingly essential. For example, a medical diagnostic system may need to interpret CT scans, extract structured information from clinical notes, and reason about temporal patterns in patient histories, while a financial forecasting system must effectively analyze and interpret plot data to make informed predictions about market trends and economic risks.
@@
Figure 1: Overview of the MEXA Architecture. Given the input task context and question, MEXA first employs an MLLM router (Sec. 3.2.2) to select the appropriate experts based on input modality and required reasoning skills. The aggregator (Sec. 3.2.3) then reasons over the outputs from the selected experts to generate the final answer.
@@ 2 Related Work
Recently, the Mixture-of-Experts (MoE) paradigm Shazeer et al. (2017); Zoph et al. (2022); Chen et al. (2023b); Zhou et al. (2022); Dai et al. (2022); Zhang et al. (2023); Chowdhury et al. (2023); Lee et al. (2024), which integrates multiple specialized parametric expert modules, has been widely adopted across a range of machine learning tasks to leverage complementary capabilities. These methods typically employ sparsely-gated mechanisms that activate only a subset of expert modules during each forward pass, improving computational efficiency and scalability. Building on this motivation, recent advances have explored leveraging a mixture of expert models (or agents(Li et al., 2024b; Wang et al., 2024; Chen et al., 2025; Li et al., 2025; Cao et al., 2025) that combine independently pre-trained models across diverse knowledge domains, moving beyond the traditional layer-wise sparse activation used in MoE-based approaches. Rather than operating as interchangeable sub-modules within a single model, such agent-based (or model-based) mixtures selectively aggregate the outputs of independently trained models through dynamic routing mechanisms to solve tasks. This design allows for more flexible and targeted knowledge utilization, promoting improved generalization across tasks and greater extensibility. However, existing approaches primarily focus on single-modality settings Chen et al. (2025) or lack support for complex reasoning tasks Li et al. (2024b); Wang et al. (2024); Li et al. (2025); Cao et al. (2025). In contrast, our method extends this line of work by scaling expert modularity to handle heterogeneous multimodal inputs and reasoning, enabling more versatile and adaptive problem-solving across domains.
@@
Real-world environments are increasingly dynamic, requiring the need for AI systems to perceive and process a broader range of modalities beyond unimodal learning. This growing demand has led to advances in models that integrate diverse signals such as images, audio, text, and 3D point clouds to improve learning and generalization. Among these, Vision-Language Models (VLMs) (Huang et al., 2023; Li et al., 2023a; Gong et al., 2023; Chen et al., 2023a) support tasks like speech recognition and audio captioning by fusing acoustic and textual information. Similarly, 2D-3D Joint Models (Li et al., 2020; Hou et al., 2021, 2023; Lei et al., 2024) combine spatial and geometric features for enhanced 3D scene understanding. However, these models are typically limited to fixed modality-task combinations and struggle to generalize or scale to novel modality inputs. This motivates the development of flexible many-modality systems (Zellers et al., 2022; Han et al., 2023; Li et al., 2023b; Girdhar et al., 2023; Liu et al., 2023b; Yu et al., 2024) that adapt to heterogeneous inputs and reasoning demands across domains. Yet, most existing systems are primarily designed to fuse multiple input modalities and learn perception and reasoning implicitly within a unified architecture. While effective for many tasks, these approaches often struggle with scalability and tend to overlook the explicit use of LLMs’ advanced reasoning capabilities. In contrast, our modular framework selects and coordinates expert models through an LLM-driven router and aggregator, directly leveraging reasoning abilities that are insufficiently exploited by unified multimodal architectures.
@@ 3 MEXA: Dynamic Multi-Modal Expert Aggregation
We first highlight the challenges faced by general-purpose multimodal reasoning systems in Sec. 3.1 and introduce our skill-specialized expert aggregation framework as a solution in Sec. 3.2. Within Sec. 3.2, we first describe the organization of a pool of expert modules and our strategy for extracting expertise-specific text representations from the specialized experts(Sec. 3.2.1). Next, we introduce an expert selection module that dynamically determines which experts to activate based on the input query, target modality, required skill, and reasoning complexity (Sec. 3.2.2). Finally, we present our expert aggregation strategy, which integrates the outputs from the selected expert to generate a final answer (Sec. 3.2.3).
@@ 3.1 General-Purpose Multimodal Reasoning
Existing general-purpose multimodal reasoning systems Han et al. (2023); Liu et al. (2023b); Yu et al. (2024) typically rely on fixed fusion architectures that process raw multimodal inputs within a shared representation space. These systems aim to implicitly learn both multimodal alignment and reasoning capabilities within a single end-to-end model. Formally, such a system can be framed as a question-answering model that takes a natural language question $Q_{\text{ques}}$ and a collection of supported modalities, such as 2D images, 3D point clouds, videos, text, and audio, denoted as =$\mathcal{M}=\{m_{1},m_{2},\dots,m_{n}\}$, where $n$ is the number of input modalities as input. The input modalities are first encoded and fused into a joint representation space, which is then processed by a monolithic reasoning model trained to reason over this fused representation. The model produces a contextually grounded answer $A$, defined as:
@@ 3.2.2 Expert Selection Module
$\begin{split}\mathcal{T}=\{x_{s}=&E_{s}(m_{s},p_{s})\ |\ E_{s}\in\texttt{Router}(\\ &Q_{\text{ques}},T_{r},\mathcal{M}),s\in\mathcal{S}\subseteq\mathcal{M}\}\\ \end{split}$ (4)
@@ 4.1 Experimental Setup
GPT-4o Hurst et al. (2024) $61.2$ $66.0$ $62.0$ 55.755.7{55.7}55.7  $69.5$ $66.9$ $51.6$ 64.864.8{64.8}64.8 $69.5$ $57.1$
Claude-3.5-Sonnet Anthropic (2024) $65.8$ 72.072.0{72.0}72.0 69.769.7{69.7}69.7 55.755.7{55.7}55.7  $66.7$ 75.075.0{75.0}75.0 56.156.1{56.1}56.1 $58.1$ 75.275.2{75.2}75.2 66.166.1{66.1}66.1
@@ 4.3 Quantitative Results
Audio-based Multimodal Reasoning. Tab. 3 presents the results of the MMAU benchmark, evaluating models on three audio QA types: Sound, Music, and Speech. Our proposed MEXA consistently outperforms all audio large language models, including GAMA, MuLLaMA, and SALAMONN, across all categories. Notably, MEXA achieves a substantial improvement of $4.1\%$, $5.9\%$, and $26.3\%$ over the best-performing Audio LLM baselines, demonstrating its strong and generalizable ability to understand diverse non-visual sensory inputs, such as audio. In contrast, existing models struggle particularly with complex music and speech QA tasks. Compared to GPT-4o with generated captions using Kim et al. (2024); Liu et al. (2024a); Radford et al. (2023), MEXA achieves comparable overall performance while maintaining a more balanced accuracy across all audio types. These results validate the effectiveness of our modular expert-based approach in addressing the diverse challenges of audio understanding.
3D Scene Understanding. Results on the SQA3D benchmark shown in Tab. 3 demonstrate the effectiveness of our method in 3D situated reasoning tasks. By integrating situated captions and general scene descriptions through our routing and aggregation system, MEXA consistently improves accuracy across all question types. Specifically, our approach achieves a $2\%$ accuracy gain over SOTA 3D-based LLMs, which rely on general-purpose encoder-fusion strategies with static unified architectures. This improvement highlights the advantage of our dynamic, modality-aware expert selection and aggregation framework in the 3D domain.
Medical QA. To demonstrate the effectiveness of our approach beyond general domains such as audio and video, we additionally evaluate MEXA on the medical video QA domain using the M3D benchmark. This benchmark includes five distinct medical QA types: Plane classification, Phase recognition, Organ identification, Abnormality detection, and Location estimation. As shown in Tab. 5, MEXA significantly outperforms strong general-purpose MLLM, CREMA, MiniCPM-o, and GPT-4o, which directly process 3D scan images as input, and observe that MEXA consistently achieves superior performance across most categories, improving average accuracy by 36.1%, 8.7% and 1.6%, respectively. These results underscore MEXA’s robust capability in medical content understanding, even without extensive domain-specific fine-tuning.
@@ 4.4 Ablation Study and Analysis
In Fig. 2, we visualize the expert selection distributions from MEXA’s expert selection module. For the Video-MMMU benchmark, multiple experts are frequently selected, aligning well with the benchmark’s inherent multimodal and multidisciplinary nature, which demands diverse skills for comprehensive video understanding. In the MMAU audio reasoning benchmark, the audio-specific experts (music, audio, and speech) are consistently activated, with their selections evenly distributed, reflecting balanced reliance on all audio modalities. For the 3D reasoning task (SQA3D), the situated 3D and general 3D scene experts are predominantly selected, with occasional inclusion of the image descriptor expert. Finally, for the medical dataset (M3D-VQA), the CT scan expert is primarily selected, complemented occasionally by the medical image expert. These observations confirm that our selection module effectively and adaptively matches experts to the specific modality requirements and reasoning contexts of each task. We provide qualitative examples of Video-MMMU and SQA3D in the Appendix.
@@ 6 Acknowledgment
This work was supported by the National Institutes of Health (NIH) under other transactions 1OT2OD038045-01, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, DARPA ECOLE Program No. HR00112390060, and NSF-AI Engage Institute DRL211263. The views, opinions, and/or findings contained in this article are those of the authors and not of the funding agency.
@@ A.2 Evaluation Datasets
MMAU (Sakshi et al., 2024) is a benchmark designed to evaluate multimodal audio understanding with curated audio clips paired with human-annotated natural language questions spanning speech, environmental sounds, and music. Evaluation is conducted on a $1$K validation set with available ground truth annotations.
SQA3D (Ma et al., 2023b) is a benchmark specifically designed for evaluating situated 3D scene understanding. It presents situations such as "You are standing beside a table", requiring models to reason about 3D spatial relationships between the described viewpoint and surrounding objects. In our evaluation, we use the SQA3D test set, which includes around $3$K human-annotated question-answer pairs.
@@
A.3 Qualitative Examples

### v2

@@ Abstract
Shoubin Yu1,thanks: Equal contribution.   Yue Zhang1,∗   Ziyang Wang1 Jaehong Yoon1,2Mohit Bansal1 1UNC Chapel Hill   2Nanyang Technological University https://github.com/Yui010206/MEXA
@@ 1 Introduction
The rapid advancement of multimodal learning has significantly improved AI systems’ ability to understand, reason, and interact with the real world, benefiting diverse tasks such as visual question answering Antol et al. (2015); Hudson and Manning (2019); Marino et al. (2019); Mathew et al. (2021); Tanaka et al. (2023); Winata et al. (2024); Wang et al. (2025); Zhang et al. (2024b); Guo et al. (2025), medical image diagnosis Lau et al. (2018); He et al. (2020); Liu et al. (2021); Azam et al. (2022); Cai et al. (2023); Bai et al. (2024), and embodied AI Brohan et al. (2023); Driess et al. (2023); Liu et al. (2024b); Durante et al. (2024); Majumdar et al. (2024); Zhang et al. (2024c). As AI systems become more integrated into complex, real-world applications, the ability to flexibly understand and reason upon heterogeneous multimodal inputs is increasingly essential. For example, a medical diagnostic system may need to interpret CT scans, extract structured information from clinical notes, and reason about temporal patterns in patient histories, while a financial forecasting system must effectively analyze and interpret plot data to make informed predictions about market trends and economic risks.
@@
Figure 1: Overview of the MEXA Architecture. Given the input task context and question, MEXA first employs an MLLM router (Sec.˜3.2.2) to select the appropriate experts based on input modality and required reasoning skills. The aggregator (Sec.˜3.2.3) then reasons over the outputs from the selected experts to generate the final answer.
@@ 2 Related Work
Recently, the Mixture-of-Experts (MoE) paradigm Shazeer et al. (2017); Zoph et al. (2022); Chen et al. (2023b); Zhou et al. (2022); Dai et al. (2022); Zhang et al. (2023); Chowdhury et al. (2023); Lee et al. (2024), which integrates multiple specialized parametric expert modules, has been widely adopted across a range of machine learning tasks to leverage complementary capabilities. These methods typically employ sparsely-gated mechanisms that activate only a subset of expert modules during each forward pass, improving computational efficiency and scalability. Building on this motivation, recent advances have explored leveraging a mixture of expert models (or agents(Li et al., 2024b; Wang et al., 2024; Chen et al., 2025; Li et al., 2025; Cao et al., 2025) that combine independently pre-trained models across diverse knowledge domains, moving beyond the traditional layer-wise sparse activation used in MoE-based approaches. Rather than operating as interchangeable sub-modules within a single model, such agent-based (or model-based) mixtures selectively aggregate the outputs of independently trained models through dynamic routing mechanisms to solve tasks. This design allows for more flexible and targeted knowledge utilization, promoting improved generalization across tasks and greater extensibility. However, existing approaches primarily focus on single-modality settings Chen et al. (2025) or lack support for complex reasoning tasks Li et al. (2024b); Wang et al. (2024); Li et al. (2025); Cao et al. (2025); Zhang et al. (2025). In contrast, our method extends this line of work by scaling expert modularity to handle heterogeneous multimodal inputs and reasoning, enabling more versatile and adaptive problem-solving across domains.
@@
Real-world environments are increasingly dynamic, requiring the need for AI systems to perceive and process a broader range of modalities beyond unimodal learning. This growing demand has led to advances in models that integrate diverse signals such as images, audio, text, and 3D point clouds to improve learning and generalization. Among these, Vision-Language Models (VLMs) (Huang et al., 2023; Li et al., 2023a; Gong et al., 2023; Chen et al., 2023a) support tasks like speech recognition and audio captioning by fusing acoustic and textual information. Similarly, 2D-3D Joint Models (Li et al., 2020; Hou et al., 2021, 2023; Lei et al., 2024) combine spatial and geometric features for enhanced 3D scene understanding. However, these models are typically limited to fixed modality-task combinations and struggle to generalize or scale to novel modality inputs. This motivates the development of flexible many-modality systems (Zellers et al., 2022; Han et al., 2023; Li et al., 2023b; Girdhar et al., 2023; Liu et al., 2023b; Yu et al., 2024) that adapt to heterogeneous inputs and reasoning demands across domains. Yet, most existing systems are primarily designed to fuse multiple input modalities and learn perception and reasoning implicitly within a unified architecture. While effective for many tasks, these approaches often struggle with scalability and tend to overlook the explicit use of LLMs’ advanced reasoning capabilities. Our modular framework selects and coordinates expert models through an LLM-driven router and aggregator, directly leveraging reasoning abilities that are insufficiently exploited by unified multimodal architectures. Another line of work, programming-based methods Lu et al. (2023); Surís et al. (2023), relies on pre-defined functions or in-context examples (few-shot prompts) and executes reasoning in a code format. In contrast, our MEXA does not require in-context examples and operates entirely in a training-free, zero-shot setting. MEXA performs deep reasoning using a Large Reasoning Model rather than code execution, allowing for free-form input rather than rigid programmatic steps. This flexibility makes MEXA more extensible and generalizable to a wider range of tasks and modalities, including audio, 3D, and medical tasks, which are beyond the scope of ViperGPT Surís et al. (2023) and Chameleon Lu et al. (2023).
@@ 3 MEXA: Dynamic Multi-Modal Expert Aggregation
We first highlight the challenges faced by general-purpose multimodal reasoning systems in Sec.˜3.1 and introduce our skill-specialized expert aggregation framework as a solution in Sec.˜3.2. Within Sec.˜3.2, we first describe the organization of a pool of expert modules and our strategy for extracting expertise-specific text representations from the specialized experts(Sec.˜3.2.1). Next, we introduce an expert selection module that dynamically determines which experts to activate based on the input query, target modality, required skill, and reasoning complexity (Sec.˜3.2.2). Finally, we present our expert aggregation strategy, which integrates the outputs from the selected expert to generate a final answer (Sec.˜3.2.3).
@@ 3.1 General-Purpose Multimodal Reasoning
Existing general-purpose multimodal reasoning systems Han et al. (2023); Liu et al. (2023b); Yu et al. (2024) typically rely on fixed fusion architectures that process raw multimodal inputs within a shared representation space. These systems aim to implicitly learn both multimodal alignment and reasoning capabilities within a single end-to-end model. Formally, such a system can be framed as a question-answering model that takes a natural language question $Q_{\text{ques}}$ and a collection of supported modalities, such as 2D images, 3D point clouds, videos, text, and audio, denoted as $\mathcal{M}=\{m_{1},m_{2},\dots,m_{n}\}$, where $n$ is the number of input modalities as input. The input modalities are first encoded and fused into a joint representation space, which is then processed by a monolithic reasoning model trained to reason over this fused representation. The model produces a contextually grounded answer $A$, defined as:
@@ 3.2.2 Expert Selection Module
[…]
@@ 4.1 Experimental Setup
GPT-4o Hurst et al. (2024) $61.2$ $66.0$ $62.0$ 55.7{55.7}  $69.5$ $66.9$ $51.6$ 64.8{64.8} $69.5$ $57.1$
Claude-3.5-Sonnet Anthropic (2024) $65.8$ 72.0{72.0} 69.7{69.7} 55.7{55.7}  $66.7$ 75.0{75.0} 56.1{56.1} $58.1$ 75.2{75.2} 66.1{66.1}
@@ 4.3 Quantitative Results
Audio-based Multimodal Reasoning. Tab.˜3 presents the results of the MMAU benchmark, evaluating models on three audio QA types: Sound, Music, and Speech. Our proposed MEXA consistently outperforms all audio large language models, including GAMA, MuLLaMA, and SALAMONN, across all categories. Notably, MEXA achieves a substantial improvement of $4.1\%$, $5.9\%$, and $26.3\%$ over the best-performing Audio LLM baselines, demonstrating its strong and generalizable ability to understand diverse non-visual sensory inputs, such as audio. In contrast, existing models struggle particularly with complex music and speech QA tasks. Compared to GPT-4o with generated captions using Kim et al. (2024); Liu et al. (2024a); Radford et al. (2023), MEXA achieves comparable overall performance while maintaining a more balanced accuracy across all audio types. These results validate the effectiveness of our modular expert-based approach in addressing the diverse challenges of audio understanding.
3D Scene Understanding. Results on the SQA3D benchmark shown in Tab.˜3 demonstrate the effectiveness of our method in 3D situated reasoning tasks. By integrating situated captions and general scene descriptions through our routing and aggregation system, MEXA consistently improves accuracy across all question types. Specifically, our approach achieves a $2\%$ accuracy gain over SOTA 3D-based LLMs, which rely on general-purpose encoder-fusion strategies with static unified architectures. This improvement highlights the advantage of our dynamic, modality-aware expert selection and aggregation framework in the 3D domain.
Medical QA. To demonstrate the effectiveness of our approach beyond general domains such as audio and video, we additionally evaluate MEXA on the medical video QA domain using the M3D benchmark. This benchmark includes five distinct medical QA types: Plane classification, Phase recognition, Organ identification, Abnormality detection, and Location estimation. As shown in Tab.˜5, MEXA significantly outperforms strong general-purpose MLLM, CREMA, MiniCPM-o, and GPT-4o, which directly process 3D scan images as input, and observe that MEXA consistently achieves superior performance across most categories, improving average accuracy by 36.1%, 8.7% and 1.6%, respectively. These results underscore MEXA’s robust capability in medical content understanding, even without extensive domain-specific fine-tuning.
@@ 4.4 Ablation Study and Analysis
In Fig.˜2, we visualize the expert selection distributions from MEXA’s expert selection module. For the Video-MMMU benchmark, multiple experts are frequently selected, aligning well with the benchmark’s inherent multimodal and multidisciplinary nature, which demands diverse skills for comprehensive video understanding. In the MMAU audio reasoning benchmark, the audio-specific experts (music, audio, and speech) are consistently activated, with their selections evenly distributed, reflecting balanced reliance on all audio modalities. For the 3D reasoning task (SQA3D), the situated 3D and general 3D scene experts are predominantly selected, with occasional inclusion of the image descriptor expert. Finally, for the medical dataset (M3D-VQA), the CT scan expert is primarily selected, complemented occasionally by the medical image expert. These observations confirm that our selection module effectively and adaptively matches experts to the specific modality requirements and reasoning contexts of each task. We provide qualitative examples of Video-MMMU and SQA3D in the Appendix.
@@ 6 Acknowledgment
We thank the reviewers and area chairs for their helpful feedback. This work was supported by the National Institutes of Health (NIH) under other transactions 1OT2OD038045-01, ARO Award W911NF2110220, ONR Grant N00014-23-1-2356, DARPA ECOLE Program No. HR00112390060, NSF-AI Engage Institute DRL211263, and a Capital One Research Award. The views, opinions, and/or findings contained in this article are those of the authors and not of the funding agency.
@@ A.2 Evaluation Datasets
MMAU (Sakshi et al., 2024) is a benchmark designed to evaluate multimodal audio understanding with curated audio clips paired with human-annotated natural language questions spanning speech, environmental sounds, and music. Evaluation is conducted on a 11K validation set with available ground truth annotations.
SQA3D (Ma et al., 2023b) is a benchmark specifically designed for evaluating situated 3D scene understanding. It presents situations such as "You are standing beside a table", requiring models to reason about 3D spatial relationships between the described viewpoint and surrounding objects. In our evaluation, we use the SQA3D test set, which includes around 33K human-annotated question-answer pairs.
@@
A.3 Running Efficiency Evaluation
In Tab.˜6, we provide detailed computational cost analysis, including: average expert skills calls and running time (router, captioner and aggregator) as listed in Table A. We observe that Video-MMMU requires, on average, more skills than other tasks. This is because video tasks contain richer multimodal information and therefore need more experts or skills to handle effectively. Additionally, our system allows a flexible trade-off between running time and accuracy by adjusting the size of the expert pool in the router prompt.
Dataset Avg. Skills Call Avg. Running Time
VideoMMMU 2.10 21.10
SQA3D 1.81 18.03
MMAU 1.10 13.31
M3D 1.10 13.81
Table 6: Average skill calls (times) and running time (s) across datasets/benchmarks.
A.4 Comparison With Other Agent Systems
In Tab.˜7, we provide extra results with the recent multimodal multi-agent system. Our MEXA achieves better performance on Video-MMMU and shows more generalizability over different tasks, while other methods are limited to the specific task.
Method Video-MMMU MMAU
VideoMultiAgent Kugo et al. (2025) $62.88$ not applicable
ReAgent-V Zhou et al. (2025) $61.11$ not applicable
MEXA $71.47$ $45.9$
Table 7: Comparison with other Agent-based works.
A.5 Extra Discussion on Expert Overlapping and Conflict
We clarify that our framework addresses redundancy and conflicts in both expert selection and aggregation. Specifically, our MLLM router dynamically selects experts based on their semantic relevance to the given query, significantly reducing unnecessary redundancy from activating all available experts. Additionally, our aggregator leverages contextual reasoning to integrate expert outputs, effectively resolving conflicts by prioritizing information consistency and relevance.
We use two examples to show how our model addresses redundancy and conflicts, respectively.
Redundancy: Consider the query: "Describe the spatial relationship between the table and the chair." Suppose both the 3D Expert and Visual Expert independently provide the response: "The chair is placed next to the table on its left side." In this scenario, our aggregator effectively integrates these overlapping outputs, boosting confidence in the identified spatial relationship ("left") between the table and chair.
Conflicts: Consider the query: "Which object is closer to the door, the chair or the table?" In this scenario, suppose two different experts provide conflicting outputs. The 3D scan expert generates: "The chair is closer to the door.", while the visual expert generates: "The table appears closer to the door." This exemplifies a conflict where two experts offer totally opposing answers. Our aggregator resolves such conflicts by analyzing the query and prioritizes the 3D scan expert, as spatial depth is more accurately captured through 3D geometry than 2D visual appearance.
A.6 Qualitative Examples